OpenStudy (anonymous):

Brennan is making a poster for the drama club’s new production. It is a regular pentagon with side lengths of 12 inches. The school wants to put up a giant replica of the poster during athletic events. If the length of each side is 8 times the original, how many times larger is the area of the replica than the area of the original?

OpenStudy (anonymous):

can you show me the steps in how you got the answer

OpenStudy (anonymous):

Original Pentagon: Sides = 5 Side Length = 12 inches Perimeter of Original Pentagon = 5*12 = 60 inches Replica Pentagon Sides = 5 Side Length = 8*Original Length = 8*12 = 96 inches Perimeter of Replica Pentagon = 5*96 = 480 inches. Perimeter of Replica Pentagon/Perimeter of Original Pentagon = 480/60 = 8. Therefore, the replica pentagon will be 8 times as larger the original pentagon.